# A Recursive Theory of Variational State Estimation: The Dynamic Programming Approach

### v1

@@ 4 On the intractability of the value functionals
The fact that the value functional recursions of theorem˜1, theorem˜2, and definition˜6 are in general intractable has hitherto not been discussed. However, $\alpha_{t}$ ought to be an unnormalized density (see corollary˜3). Therefore, some care needs to taken to ensure that $\alpha_{t}$ remains normalizable throughout recursion. Here this issue shall not be examined in exhaustive detail, there is simply too many possible approaches that will have to be explored elsewhere. However, there is a simple approach approach will be offered. Namely, by definition˜5, the normalization constant of $\alpha_{t}$ has a lower bound according to
$\begin{split}\log\int\alpha_{t}(x_{t})\operatorname{d\!}x_{t}&\geq\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}\Big\}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t}\\ &=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}.\end{split}$ (49)
@@
$\underset{q_{t-1}\in\mathcal{Q}_{t-1},\,q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}{\max}\,\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\hat{q}_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t\mid t-1}q_{t-1}}\Big\}q_{t-1\mid t}q_{t}\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t},$ (51)
@@ 6 Case study: Jump Gauss–Markov systems
Consider the following Markov chain
@@
The backward and forward value functionals are for $0\leq t\leq T$ respectively given by
@@ 6.2 Simulation study: an autoregressive staircase
There are various recursive approaches to computing filtering densities and smoothing marginals in jump Gauss–Markov systems (BarShalom2004, Helmick1995, Balenzuela2022). However, to the authors knowledge, none that can compute a posterior over the entire path $(x_{0\mathrel{\mathop{\ordinarycolon}}T},z_{0\mathrel{\mathop{\ordinarycolon}}T})$. In any case, the posterior obtained from proposition˜2 is optimal in relative entropy sense with respect to the chosen variational family. Thus instead the sub-optimal variational filter (VJGM(0)) is compared to an interacting multiple model filter with $M$ components (IMM) (BarShalom2004), to ensure the notion of sub-optimal vairational filtering is reasonable. Additionally, (VJGM(0)) is used to initialize the fixed-point iteration of proposition˜2, which runs for 10 iterations (VJGM(10)).  [^6] It seems to always converge in less than 10 iterations in this example. The posteriors obtained from (VJGM(0)) and (VJGM(10)) are then compared. The parameters are set to $\mu_{0}(z)=z-1$, $\Sigma_{0}(z)=0.5^{2}$, $\phi_{0}=0.5$, $M=4$, $p=0.9$, $R=1$, and the number of time points is set to $T=2^{9}+1$. A realization of the system is shown in fig.˜1.
@@ 7 Conclusion
In this article, a complete set of variational analogue to Bayesian state estimation theory was established. In particular, the value functional associated with the backward method, $\beta_{t}$, was shown to be a lower bound on the state likelihood $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$, and the value functional associated with the forward method, $\alpha_{t}$, was shown to be a lower bound on the unnormalized filtering density $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. Like the Bayesian state estimation recursions, the variational state estimation recursion are intractable in general. However, it has been demonstrated that they are tractable more often than their Bayesian counterparts. Additionally, a formal notion of sub-optimal variational filtering has been put forth, which ensures a time-complexity of $O(T)$. Notably this notion of variational filtering differs from that of Campbell2021. The methodology has been empirically demonstrated to be sound in a jump Gauss–Markov system with a factored approximation. There are several issues and possibilities that can be pursued in future work. From a methodological point of view, the two most pressing ones appear to be:
@@ Appendix A More details on jump Gauss–Markov processes
In this section, some additonal details are provided for variational state estimation in jump Gauss–Markov systems. More specieifcally formulae are provided for conveniently organizing the implementation. The proofs of these results reduce to rote algebra and are therefore omitted.

### v2

@@ 4 On the intractability of the value functionals
The fact that the value functional recursions of theorem˜1, theorem˜2, and definition˜6 are in general intractable has hitherto not been discussed. However, $\alpha_{t}$ ought to be an unnormalized density (see corollary˜3). Therefore, some care needs to taken to ensure that $\alpha_{t}$ remains normalizable throughout recursion. Here this issue shall not be examined in exhaustive detail, there is simply too many possible approaches that will have to be explored elsewhere. However, there is a simple approach approach will be offered. Namely, by definition˜6, the normalization constant of $\alpha_{t}$ has a lower bound according to
$\begin{split}\log\int\alpha_{t}(x_{t})\operatorname{d\!}x_{t}&\geq\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t}\\ &=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}.\end{split}$ (50)
@@
$\underset{q_{t-1}\in\mathcal{Q}_{t-1},\,q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}{\max}\,\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}q_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t\mid t-1}q_{t-1}}\Big\}q_{t-1\mid t}q_{t}\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t},$ (52)
@@ 5 Related work
A few days later after the first pre-print of the present manuscript was released, another approach appeared by (Abdulsamad2025). Where they study iterative recursive variational Gaussian smoothing under entropic regularization using a Lagrange multiplier approach rather than dynamic programming. Also, rather than imposing a constraint class $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ explicitly they “close” the value functional recursions by a variety of quadratic approximations.
@@ 6 Case study: Jump Gauss–Markov systems
In this section the variational forward method is studied in jump Gauss–Markov systems under a certain factored approximation, which leads to an alternate coordinate ascent method for computing the transition kernels. More specifically, consider the following Markov chain
@@
The forward value functional is for $0\leq t\leq T$ respectively given by
@@ 6.2 Simulation study: an autoregressive staircase
There are various recursive approaches to computing filtering densities and smoothing marginals in jump Gauss–Markov systems (BarShalom2004, Helmick1995, Balenzuela2022). However, to the authors knowledge, none that can recursively compute a posterior over the entire path $(x_{0\mathrel{\mathop{\ordinarycolon}}T},z_{0\mathrel{\mathop{\ordinarycolon}}T})$. In any case, the posterior obtained from proposition˜2 is optimal in relative entropy sense with respect to the chosen variational family. Thus instead the sub-optimal variational filter (VJGM(0)) is compared to an interacting multiple model filter with $M$ components (IMM) (BarShalom2004), to ensure the notion of sub-optimal vairational filtering is reasonable. Additionally, (VJGM(0)) is used to initialize the fixed-point iteration of proposition˜2, which runs for 10 iterations (VJGM(10)).  [^5] It seems to always converge in less than 10 iterations in this example. The posteriors obtained from (VJGM(0)) and (VJGM(10)) are then compared. The parameters are set to $\mu_{0}(z)=z-1$, $\Sigma_{0}(z)=0.5^{2}$, $\phi_{0}=0.5$, $M=4$, $p=0.9$, $R=1$, and the number of time points is set to $T=2^{9}+1$. A realization of the system is shown in fig.˜1.
@@
7 Case study: log-polynomial systems
In this section, variational filtering is studied when the problem is log-polynomial of order, say $\nu$. The model is then given by
$\log\bar{\pi}_{0}^{0}(x_{0})$ $=\log\mathcal{N}(x_{0};\mu_{0},\Sigma_{0})+\sum_{m=0}^{\nu}\ell_{0}^{(m)}x_{0}^{m}$ (78a)
$\log\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})$ $=\log\mathcal{N}(x_{t};a_{t\mid t-1}(x_{t-1}),Q_{t\mid t-1})+\sum_{m=0}^{\nu}\ell_{t}^{(m)}x_{t}^{m}$ (78b)
and the state is thus a scalar. Note that by defining $p_{m}(x)=x^{m}$ the likelihoods can be written as inner products according to
$\log h_{t\mid t}(x_{t})=\sum_{m=0}^{\nu}\ell_{t}^{(m)}x_{t}^{m}=\langle\ell_{t},p(x_{t})\rangle.$ (79)
The aim of this section is to establish the forward value functional recursion when $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ is Gauss–Markov and thus the following theorem.
Theorem 7.
Let the variational family $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ be Gauss–Markov then the forward value functional associated with a log-polynomial system of order $\nu\geq 2$ is given by
$\log\alpha_{t}(x_{t})=\sum_{m=0}^{\nu}\lambda_{t}^{(m)}x_{t}^{m}=\langle\lambda_{t},p(x_{t})\rangle.$ (80)
Theorem˜7 readily gives the filter update, namely
$\begin{split}\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}&=\int\langle\lambda_{t},p(x_{t})\rangle q_{t}(x_{t})\operatorname{d\!}x_{t}-\int\log\{q_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$ (81)
Applying lemma˜8 from appendix˜B to the first term then gives the following result.
Proposition 7 (Variational filter update).
Let $q_{t}$ ge Gaussian with mean and covariance $\mu_{t}$ and $\Sigma_{t}$, respectively, then
$\begin{split}\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}&=\langle\lambda_{t},A_{1}(\Sigma_{t})p(\mu_{t})\rangle+\frac{1}{2}(\log\{2\pi\Sigma_{t}\}+1),\end{split}$ (82)
where $A_{1}$ is defined in lemma˜8. Hence the variational filter mean and covariance satisfies
$\mu_{t}^{0\mathrel{\mathop{\ordinarycolon}}t},\,\Sigma_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\underset{\mu_{t},\Sigma_{t}}{\arg\,\max}\,\big\{\langle\lambda_{t},A_{1}(\Sigma_{t})p(\mu_{t})\rangle+\frac{1}{2}(\log\{2\pi\Sigma_{t}\}+1)\big\}.$ (83)
It now remains to establish theorem˜7 and a recursion for the value functional. Note that theorem˜7 hold at $t=0$ by the following identification
$\lambda_{0}^{(0)}$ $=-\frac{1}{2}\log\{2\pi\Sigma_{0}\}-\frac{\mu_{0}^{2}}{2\Sigma_{0}}+\ell_{0}^{(0)}$ (84a)
$\lambda_{0}^{(1)}$ $=\frac{\mu_{0}}{\Sigma_{0}}+\ell_{0}^{(1)}$ (84b)
$\lambda_{0}^{(2)}$ $=-\frac{1}{2\Sigma_{0}}+\ell_{0}^{(2)}$ (84c)
$\lambda_{0}^{(m)}$ $=\ell_{0}^{(m)},\quad 2\leq m.$ (84d)
It then remains to establish the induction step. The value functional recursion is
[…]
where (79) was used for the first term and $c_{t}$ is defiend by lemma˜9 in appendix˜B. Additionally, by lemma˜8 in appendix˜B, there are matrices $A_{1}(Q_{t-1\mid t})$ and $A_{2}(\Phi_{t-1,t},u_{t-1,t})$ such that
$\begin{split}\int\langle\lambda_{t-1},p(x_{t-1})\rangle q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}&=\langle\lambda_{t-1},A_{1}(Q_{t-1\mid t})p(a_{t-1\mid t}(x_{t}))\rangle\\ &=\langle\lambda_{t-1},A_{1}(Q_{t-1\mid t})A_{2}(\Phi_{t-1,t},u_{t-1,t})p(x_{t})\rangle\end{split}$ (86)
Theorem˜7 has then been established along with the following value functional recursion.
Proposition 8 (Forward value functional update).
Assume theorem˜7 hold at t1t-1 then it also hold at $t$. Additionally the parameter $\lambda_{t}$ can be obtained from $\lambda_{t-1}$ according to
$\lambda_{t}=A_{2}^{*}(\Phi_{t-1,t},u_{t-1,t})A_{1}^{*}(Q_{t-1\mid t})\lambda_{t-1}+c_{t}(\Phi_{t-1,t},u_{t-1,t},Q_{t-1\mid t})+\ell_{t}$ (87)
where the matrices $A_{1}$ and $A_{2}$ are defined in lemma˜8, and $c_{t}$ is defined in lemma˜9.
7.1 Simulation study: The cubic sensor problem
Let $a_{t\mid t-1}(x)=\phi_{0}x+(1-\phi_{0})\mu_{0}$ and $c_{t\mid t}(x)=\beta x^{3}$ and consider the cubic sensor system given by (Katayama2013)
$x_{0}$ $\sim\mathcal{N}(\mu_{0},\Sigma_{0})$ (88a)
$x_{t}\mid x_{t-1}$ $\sim\mathcal{N}(x_{t};a_{t\mid t-1}(x_{t-1}),Q_{t\mid t-1})$ (88b)
$y_{t}\mid x_{t}$ $\sim\mathcal{N}(y_{t};c_{t\mid t}(x_{t}),R).$ (88c)
It is clearly a log-polynomial system of order $6$ with the likelihoods given by
$\begin{split}\log h_{t\mid t}(x_{t})&=-\frac{1}{2}\Big(\log\{2\pi R\}+\frac{y_{t}^{2}}{2R}\Big)+\frac{y_{t}\beta}{R}x_{t}^{3}-\frac{\beta^{2}}{2R}x_{t}^{6}\\ &=\sum_{m=0}^{6}\ell_{t}^{(m)}x_{t}^{m}.\end{split}$ (89)
It is interesting to note that this sytsem has two regimes. Namely, $(i)$ $\mathinner{\!\left\lvert x_{t}\right\rvert}<1$ where the signal-to-noise ratio is small and $(ii)$ $\mathinner{\!\left\lvert x_{t}\right\rvert}\geq 1$ where the signal-to-noise-ratio is high. In particular, if $\mu_{0}=0$ then $x_{t}$ is a zero mean stationary process, which together with (89) means that the posterior $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ has a mode at the zero-path, $x_{t}=0,\quad t\geq 0$.
The parameters are set to $\Sigma_{0}=0.36$, $\phi_{0}=0.95$, $Q_{t\mid t-1}=(1-\mathinner{\!\left\lvert\phi_{0}\right\rvert}^{2})\Sigma_{0}$, $\beta=1$, $R=1$, and the mean parameter is varied according to $\mu_{0}\in\{0,0.5,1.5\}$. Thus when $\mu_{0}=0$, the state has a 10% chance of being greater than unity in magnitude. The system is for $T=2^{13}+1$ time points for each configuration.
The sub-optimal variational filter of definition˜6 is implemented and the optimization is carried out with natural gradient descent using a step-size of $1\text{\times}{10}^{-1}$ and 20 iterations per filter step. The sub-optimal filter is compared to an iterated extended Kalman filter (Bell1993), which also uses 20 iterations. The two alternatives are referred to as VF(20) and EKF(20), respectively.
Figure 4: Simulation results for the cubic sensor experiment. The rop row shows the realization of the observation sequence for each selection of $\mu_{0}$ and the bottom row shows the corresponding realization of the state sequence. The filter estimates are illustrated by a two standard deviation band around their estimates.
The results of the three simulations are shown in fig.˜4. It can be seen that EKF(20) fails to track the state completely in the case when the stationary mean of the state is zero (it gets stuck at the mode at zero). However, as the stationary mean moves away from zero VF(20) and EKF(20) become increasingly similar in performance. This is also corroborated by the sequence of root mean-square errors, which is shown in table˜1. The sum of $\chi^{2}$-statistics is shown in table˜2 from which a similar conclusion can be drawn but also that both filters are quite severely overconfident on average. In fact, the overconfidence gets worse the more time the state stays in the high signal-to-noise ratio regime.
Table 1: Root mean-square errors for VF(20) and EKF(20) in the cubic sensor problem.
It can be seen that VF(20) handles the cases when the state is seldom in the high signal-to-noise ratio regime better than EKF(20).
$\mu_{0}$ 0 0.4 0.8
VF(20) 0.411192 0.380229 0.33623
EKF(20) 0.614548 0.527413 0.410902
Table 2: Sum of $\chi^{2}$-statistics for VF(20) and EKF(20) in the cubic sensor problem. If a Gaussian filter is appropriate this sum should with high probability be $T=2^{13}+1=8193$
$\mu_{0}$ 0 0.4 0.8
VF(20) 11466.5 13863.3 20122.9
EKF(20) 8595.12 27356.0 19681.2
@@ 8 Conclusion
In this article, a complete set of variational analogue to Bayesian state estimation theory was established. In particular, the value functional associated with the backward method, $\beta_{t}$, was shown to be a lower bound on the state likelihood $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$, and the value functional associated with the forward method, $\alpha_{t}$, was shown to be a lower bound on the unnormalized filtering density $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. Like the Bayesian state estimation recursions, the variational state estimation recursion are intractable in general. However, it has been demonstrated that they are tractable more often than their Bayesian counterparts. Additionally, a formal notion of sub-optimal variational filtering has been put forth, which ensures a time-complexity of $O(T)$. Notably this notion of variational filtering differs from that of Campbell2021. The methodology has been empirically demonstrated to be sound in a jump Gauss–Markov system with a factored approximation and in a Gauss–Markov model with log-polynomial likelihoods under a Gauss–Markov constraint. There are several issues and possibilities that can be pursued in future work. From a methodological point of view, the two most pressing issues appear to be:
@@
A possibility of resolving the first issue is to approximate intergrals with respect to the time-marginals with Monte Carlo averages. As the value functional recursions of theorems˜3 and 4 are given in the weak sense, an immediate approach would be to resort to some Galerkin style approximation with an appropriate selection of subspace. However, these possibilities will have to explored in future work.
@@ Appendix A More details on jump Gauss–Markov processes
In this section, some additonal details are provided for variational state estimation in jump Gauss–Markov systems. More specifically formulae are provided for conveniently organizing the implementation. The proofs of these results reduce to rote algebra and are therefore omitted.
@@
Appendix B More details on log-polynomial systems
In this section additional details are givne for log-polynomial systems. Some results rely on the classic formula for central moments of univariate Gaussian distributions, which is stated in lemma˜7.
Lemma 7 (Gaussian central moments).
Let $q(x)=\mathcal{N}(x;\mu,\Sigma)$ then the central moments are given by
$\int(x-\mu)^{p}q(x)\operatorname{d\!}x=\begin{cases}0,\quad p\text{ odd}\\ \Sigma^{p/2}(p-1)!!,\quad p\text{ even}\end{cases},$ (95)
where !!!! denotes the double factorial.
Lemma 8.
Let $\kappa(x\mid z)=\mathcal{N}(x;a(z),Q)$ with conditional mean function $a(x)=\Phi x+u$ and define the matrix $A_{1}(Q)$ by
$\{A_{1}(Q)\}_{n,m}$ $=\begin{cases}{n\choose m}Q^{\frac{n-m}{2}}(n-m-1)!!,\quad n-m\text{ even}\\ 0,\quad n-m\text{ odd or }m>n\end{cases}$ (96a)
$\{A_{2}(\Phi,u)\}_{n,m}$ $={n\choose m}\Phi^{m}u^{n-m}.$ (96b)
Then the following hold
$\int p(x)\kappa(x\mid z)\operatorname{d\!}x=A_{1}(Q)p(a(z))=A_{1}(Q)A_{2}(\Phi,u)p(z)$ (97)
Lemma 9.
Let $\kappa_{1}(x\mid z)=\mathcal{N}(x;a_{1}(z),Q_{1})$ and $\kappa_{2}(z\mid x)=\mathcal{N}(z;a_{2}(x),Q_{2})$ with $a_{i}$ being an affine map with slope and intercept given by $\Phi_{i}$ and $u_{i}$, respectively, for $i=1,2$. Then the negative cross-entropy of $\kappa_{2}$ relative to $\kappa_{1}$ is given by
$\begin{split}\int\log\Big\{\frac{\kappa_{2}(z\mid x)}{\kappa_{1}(x\mid z)}\Big\}\kappa_{1}(x\mid z)\operatorname{d\!}x&=-\frac{1}{2}\Big\{\log\mathinner{\!\left\lvert Q_{2}\right\rvert}-\log\mathinner{\!\left\lvert Q_{1}\right\rvert}+\operatorname{tr}(Q_{2}^{-1}\Phi_{2}Q_{1}\Phi_{2}^{*})-\operatorname{tr}(Q_{1}^{-1}Q_{1})\Big\}\\ &\quad-\frac{1}{2}\big\langle z-a_{2}\circ a_{1}(z),Q_{2}^{-1},(z-a_{2}\circ a_{1}(z))\big\rangle.\end{split}$
Additionally, in the scalar case this evaluates to a polynomial of degree two with coefficents given by
$c^{(0)}(\Phi_{1},u_{1},Q_{1})$ $=-\frac{1}{2}\Big\{\log Q_{2}-\log Q_{1}+\frac{\mathinner{\!\left\lvert\Phi_{2}\right\rvert}^{2}Q_{1}}{Q_{2}}-1+\frac{\mathinner{\!\left\lvert a_{2}(u_{1})\right\rvert}^{2}}{Q_{2}}\Big\}$ (98a)
$c^{(1)}(\Phi_{1},u_{1},Q_{1})$ $=\frac{(1-\Phi_{2}\Phi_{1})a_{2}(u_{1})}{Q_{2}}$ (98b)
$c^{(2)}(\Phi_{1},u_{1},Q_{1})$ $=-\frac{\mathinner{\!\left\lvert 1-\Phi_{2}\Phi_{1}\right\rvert}^{2}}{2Q_{2}}$ (98c)